OpenStudy (anonymous):

Please help!! Medal+fan! In a survey 3 out of 7 people named Blue as their favorite color. 1 out of 6 named red. If 1092 people were included in the survey how many named neither blue or red as their favorite color? A. 650 people B.468 people C.442 people D. 806 people

OpenStudy (boldjon):

Since 3/7 have favorite blue and 1/6 have favorite red, fraction with favorite: 3/7 + 1/6 = 25/42 Those with neither blue non red would be: 1 - 25/42 = 17/42 The actual number of people will be : 17/42 * 1092 = 442

OpenStudy (anonymous):

Thank you so much! @boldjon

OpenStudy (boldjon):

ur welcome